# LUQ: Layerwise Ultra-Low Bit Quantization for Multimodal Large Language Models

### v1

@@ 3.1 Setup and Notation
Let $g_{\phi}$ and $f_{\theta}$ denote the vision encoder and language model of the multimodal LLM respectively, parameterized by weights $\phi,\theta$. Any given image-text input pair $(\mathbf{I},\mathbf{T})$ is converted to a sequence of multimodal tokens denoted by $\mathbf{x}$ using the model tokenizer and $g_{\phi}$. The input sequence $\mathbf{x}$, with length NN multimodal tokens is then passed through $f_{\theta}$, which is a series of $L$ transformer layers, where each layer $f_{\theta_{i}}$ ($i\in\{1,...,L\}$) transforms its input representations:
@@ 3.2 Activation Entropy Estimation
Given the set of layerwise intermediate activations $X_{i}$, we first convert all elements $x\in X_{i}$ with dimensions $N\times d$ into $X^{{}^{\prime}}_{i}$, a set of $(\mathcal{N}\times N)$ $d$-dimensional tokens, where NN is the sequence length and $d$ is the hidden dimension. We then estimate $H_{i}$, the entropy of the output distribution, by first estimating the distribution of $X^{{}^{\prime}}_{i}$ using cluster-based discretization. Specifically, we use K-means clustering to partition the $(\mathcal{N}\times N)$ token representations into $K$ distinct clusters.
@@ 3.4 Effect of Calibration Data Selection
We propose to instead use Mixed Modal Calibration, by explicitly sampling both text and visual tokens. Given a calibration budget of NN tokens, we construct our calibration set $\mathcal{D}_{\text{cal}}$ by sampling both text tokens $\mathcal{T}$ and multimodal (image) tokens $\mathcal{M}$ from the training distribution:
@@
Datasets $\rightarrow$ Methods $\downarrow$ Avg. Bit width MME Per. MME Cog. $M$ Bench Text VQA VQAv2 GQA POPE Chart QA Doc QA Math Vista LLaVA-1.5 7B Backbone FP16 (Baseline) 16 1510 350 63.4 58.2 78.5 62.0 83.2 - - 23.6 GPTQ 4 1450 347 58.2 56.8 76.3 61.4 76.0 - - 20.1 AWQ 4 1456 349 59.8 56.7 76.6 61.5 76.7 - - 20.6 GPTQ 3 1346 273 31.2 54.1 73.5 58.8 70.5 - - 16.4 GPTQ* 2 0 0 0.0 0 0 0.0 0.0 - - 0.0 AWQ* 2 0 0 0.0 0 0 0.0 0.0 - - 0.0 BiLLM 1.08 561 39 7.4 15.6 37.2 22.7 25.5 - - 3.5 LUQ 16 layer (Ours) 2.54 1365 257 46.7 53.4 74.9 58.2 74.5 - - 18.7 Qwen 2.5 VL Backbone FP16 (Baseline) 16 1695 640 82.6 84.9 83.5 60.5 86.1 87.3 95.7 68.2 GPTQ 4 1638 610 80.2 84.2 82.6 60.1 84.8 84.1 93.4 44.8 AWQ 4 1645 620 80.9 84.6 82.7 60.5 85.6 84.5 93.5 46.1 GPTQ 3 319 131 34.7 79.5 81.5 53.4 82.9 61.0 89.2 21 GPTQ* 2 0.0 0.0 0.0 0 0.0 0.0 0.0 0.0 0.0 0.0 AWQ* 2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 BiLLM 1.08 638 42 9.7 26.3 39.5 4.3 70.7 3.7 20.3 15.1 LUQ 12 layer (Ours) 2.75 1640 600 63.7 81.9 79.7 52.9 84.7 68.6 90.5 41.7

### v2

@@ 3.1 Setup and Notation
Let $g_{\phi}$ and $f_{\theta}$ denote the vision encoder and language model of the multimodal LLM respectively, parameterized by weights $\phi,\theta$. Any given image-text input pair $(\mathbf{I},\mathbf{T})$ is converted to a sequence of multimodal tokens denoted by $\mathbf{x}$ using the model tokenizer and $g_{\phi}$. The input sequence $\mathbf{x}$, with length $N$ multimodal tokens is then passed through $f_{\theta}$, which is a series of $L$ transformer layers, where each layer $f_{\theta_{i}}$ ($i\in\{1,...,L\}$) transforms its input representations:
@@ 3.2 Activation Entropy Estimation
Given the set of layerwise intermediate activations $X_{i}$, we first convert all elements $x\in X_{i}$ with dimensions $N\times d$ into $X^{{}^{\prime}}_{i}$, a set of $(\mathcal{N}\times N)$ $d$-dimensional tokens, where $N$ is the sequence length and $d$ is the hidden dimension. We then estimate $H_{i}$, the entropy of the output distribution, by first estimating the distribution of $X^{{}^{\prime}}_{i}$ using cluster-based discretization. Specifically, we use K-means clustering to partition the $(\mathcal{N}\times N)$ token representations into $K$ distinct clusters.
@@ 3.4 Effect of Calibration Data Selection
We propose to instead use Mixed Modal Calibration, by explicitly sampling both text and visual tokens. Given a calibration budget of $N$ tokens, we construct our calibration set $\mathcal{D}_{\text{cal}}$ by sampling both text tokens $\mathcal{T}$ and multimodal (image) tokens $\mathcal{M}$ from the training distribution:
@@
Datasets $\rightarrow$ Methods $\downarrow$ Avg. Bit width MME Per. MME Cog. MM Bench Text VQA VQAv2 GQA POPE Chart QA Doc QA Math Vista LLaVA-1.5 7B Backbone FP16 (Baseline) 16 1510 350 63.4 58.2 78.5 62.0 83.2 - - 23.6 GPTQ 4 1450 347 58.2 56.8 76.3 61.4 76.0 - - 20.1 AWQ 4 1456 349 59.8 56.7 76.6 61.5 76.7 - - 20.6 GPTQ 3 1346 273 31.2 54.1 73.5 58.8 70.5 - - 16.4 GPTQ* 2 0 0 0.0 0 0 0.0 0.0 - - 0.0 AWQ* 2 0 0 0.0 0 0 0.0 0.0 - - 0.0 BiLLM 1.08 561 39 7.4 15.6 37.2 22.7 25.5 - - 3.5 LUQ 16 layer (Ours) 2.54 1365 257 46.7 53.4 74.9 58.2 74.5 - - 18.7 Qwen 2.5 VL Backbone FP16 (Baseline) 16 1695 640 82.6 84.9 83.5 60.5 86.1 87.3 95.7 68.2 GPTQ 4 1638 610 80.2 84.2 82.6 60.1 84.8 84.1 93.4 44.8 AWQ 4 1645 620 80.9 84.6 82.7 60.5 85.6 84.5 93.5 46.1 GPTQ 3 319 131 34.7 79.5 81.5 53.4 82.9 61.0 89.2 21 GPTQ* 2 0.0 0.0 0.0 0 0.0 0.0 0.0 0.0 0.0 0.0 AWQ* 2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 BiLLM 1.08 638 42 9.7 26.3 39.5 4.3 70.7 3.7 20.3 15.1 LUQ 12 layer (Ours) 2.75 1640 600 63.7 81.9 79.7 52.9 84.7 68.6 90.5 41.7